5. Given an upper body weight of 61.2 kg and a distance l of 15 cm from the sacrum to each joint, calculate the torque at both the left and right hip joints.
Answer #1

mass = m = 61.2 kg

weight = F = m g = 61.2 * 9.8 = 599.76 N

R = l = 15 cm = 0.15 m

torque about left end = t = r * F = - 0.15 * 599.76 =- 89.96 N - m ( clock wise)

about right end = t = 89.96 N-m ( anti clock wise )